verify that sin(a) + 1 sin (2) - = S: (a)-1 all stops Show and label
Answer #1

Sin²(2)-1 - Sin(27-1 = siulalt 1 L.His Sin²(Q)-(15 Sin(27-1 (22-62 = (a-6)(a+b)/ (sin (2) + 1) (sin(a)al) (singen-1) = (sin(+